OpenStudy (anonymous):

3. Solve: 2x^2 – 18x + 28 = 0

OpenStudy (anonymous):

use quadratic formula..

OpenStudy (anonymous):

I dont know how.

OpenStudy (anonymous):

\[-b \pm \sqrt{b ^{2}-4ac}/2a\]

OpenStudy (anonymous):

take 2 common from ur equation: \[2x ^{2}-18x+28\] 2(x^2-9x+14) so ur a=1; b=-9; c=14

OpenStudy (anonymous):

\[-(-9)\pm \sqrt{(-9)^{2}-4(1)(14)}/2(1)\]

OpenStudy (anonymous):